You combine 0.75 moles formate and 0.85 moles formic acid to make a buffer solution. The Ka of formic acid is 1.8x104 what is the pH of the solution? Question 2 With the same initial solution of 0.75 moles formate and 0.85 moles formic acid to make a buffer solution. The Ka of formic acid is 1.8x104. You add 8g sodium hydroxide to this solution, what is the new pH? Question 3 With the same initial solution of 0.75 moles formate and 0.85 moles formic acid to make a buffer solution. The Ka of formic acid is 1.8x10-4 You add 16g hydrobromic acid to this solution, what is the new pH? Question 4 With the same initial solution of 0.75 moles formate and 0.85 moles formic acid to make a buffer solution. The Ka of formic acid is 1.8x10-4 You add 60.7g hydrobromic acid to this solution, what is the new pH?
Answer #1

Ans 1 assume 1 L of solution

[Acid] = 0.85 M;

[salt] = 0.75 M

pKa = – log{(Ka) / M} = – log (1.8 x 10-4) = 3.75

pH = pKa + log[salt] / [acid]) = 3.75 + log(0.75M / 0.85M)

pH = 3.75 + log(0.88) = 3.75 - 0.056 = 3.69

Ans 2 -

assume 1 L of solution

Add 8 g NaOH

Moles of NaOH added = 8/40 = 0.2 moles

The concentration will change

[Acid] = 0.85 - 0.2 = 0.65 M;

[salt]  = 0.75 + 0.2 = 0.95 M

pKa = – log{(Ka) / M} = – log (1.8 x 10-4) = 3.75

pH = pKa + log[salt] / [acid]) = 3.75 + log(0.95M / 0.65M)

pH = 3.75 + log(1.46) = 3.75 + 0.16 = 3.91

Ans 3

assume 1 L of solution

Add 16 g HBr

Moles of HBr added = 16/80.91 = 0.2 moles

The concentration will change

[Acid] = 0.85 + 0.2 = 1.05 M;

[salt]  = 0.75 - 0.2 = 0.55 M

pKa = – log{(Ka) / M} = – log (1.8 x 10-4) = 3.75

pH = pKa + log[salt] / [acid]) = 3.75 + log(0.55M / 1.05M)

pH = 3.75 + log(0.52) = 3.75 - 0.28 = 3.47

Ans 4

assume 1 L of solution

Add 60.7 g HBr

Moles of HBr added = 60.7/80.91 = 0.75 moles

The concentration will change

[Acid] = 0.85 + 0.75 = 1.60 M;

[salt]  = 0.75 - 0.75 = 0 M

pKa = – log{(Ka) / M} = – log (1.8 x 10-4) = 3.75

pH = pKa + log[salt] / [acid]) = 3.75 + log(0M / 1.60M)

pH = 3.75 + log(0) = 3.75 = 3.75